During the past year, you had a portfolio that contained U.S. government T-bills, long-term government bonds, and common stocks. The rates of return on each of them were as follows:

U.S. government T-bills 3.40 %

U.S. government long-term bonds 4.70

U.S. common stocks 6.20

During the year, the consumer price index, which measures the rate of inflation, went from 100 to 114 (1982 – 1984 = 100).

Compute the rate of inflation during this year. Round your answer to one decimal place. %

Compute the real rates of return on each of the investments in your portfolio based on the inflation rate. Use a minus sign to enter negative values, if any. Do not round intermediate calculations. Round your answers to two decimal places.

Real rate of return

U.S. government T-bills %:

U.S. government long-term bonds %:

U.S. common stocks %:
Answer #1

inflation = (ending CPI - starting CPI) / starting CPI = (114 - 100) / 114 = 14.0%

real rate of return = [(1 + nominal return) / (1 + inflation)] - 1

U.S. government T-bills = [(1 + 3.40%) / (1 + 14.0%)] - 1 = -0.0930, or -9.30%

U.S. government long-term bonds = [(1 + 4.70%) / (1 + 14.0%)] - 1 = -0.0816, or -8.16%

U.S. common stocks = [(1 + 6.20%) / (1 + 14.0%)] - 1 = -0.0684, or -6.84%